100g ice at 0 o C is mixed with 10g steam at 100 o C. Find the final temperature and composition.

100g ice at 0oC is mixed with 10g steam at 100oC. Find the final temperature and composition.

Grade:12

2 Answers

Rr
6 years ago
Abid
6 years ago
Latent heat of ice =m×l 100×80=8000Latent heat of steam =m×l =10×540 L.H of ice is greater than steam Specific heat of steam = mct =10×1×100=10005400+1000=6400 for steamCompositionQ=mlm=q/lm=6400/80m=80100-80=20g of ice 10+80=90g water